# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v4

@@ 5.1.3 Limits of Knowledge and Reasoning
Second, we prove that hallucinations are inevitable even when knowledge is subadditive. In that setting, by Theorem 4, reasoning does not create new information. It is a transformation that reveals information already present in any knowledge representation. Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. However, we should take care of a paradox those claims apparently bring into existence.
@@
To resolve it, we need the emergence operator $\mathcal{E}_{C}(X)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time). Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X)\leq\mu_{C}(\mathcal{E}_{C}(X))$. That is possible only with organized information providing knowledge for reasoning. These are precisely the statements of Theorem 1 and Theorem 4.
@@
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t))$ (11)
@@
$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)).$ (12)
@@
$\theta(1)$ $=\mathcal{E}_{C}(\theta(0))=[1,1,1,1,1]\text{ (state after reasoning)},$
$y(0)$ $=\mu_{C}(\theta(0))=3\text{ bits},$
$y(1)$ $=\mu_{C}(\theta(1))=5\text{ bits}.$
We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)))>\mu_{C}(\theta(0))$, i.e., practical information gain in the course of reasoning. In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0))$
@@ 5.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses with their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions). We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths. The following formal definition encapsulates those ideas.
@@
For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Indeed, that limits are given by the Goedel’s Theorems. Therefore, in practice that inaccessibility of the complete truth validation function is one of the reasons why hallucination cannot be eliminated entirely. High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth. Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
@@
A query $q\in\mathcal{Q}$ is called non-trivial if there exist at least two agents with information subsets $\theta$ and $\theta^{\prime}$ such that:
$K_{M}\cap T(q)\cap\bigl{[}K(\theta)\triangle K(\theta^{\prime})\bigr{]}\neq\varnothing.$ (14)
Intuitively, query is non-trivial when at least two agents disagree on some fact that is both encoded in the and relevant to answering that query. For trivial queries impossibility results do not hold.
@@ 5.4 Hallucination Cost
If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$. Fabricated or incomplete facts generate positive cost $J(r,q)>0$.
@@ 5.5.2 Semantic Information Conservation
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (19)
@@
Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$. Then, let us assume that the knowledge in the query-only response meets the following condition:
@@
$K_{i}\perp_{C}K_{j}\text{ for all }i\neq j\quad\text{and}\quad K_{i}\perp_{C}K_{0}\text{ for all }i.$ (22)
@@
$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}\bigr{)}.$ (23)
@@
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right),$ (24)
@@
$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$ (25)
If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$, then the response is creatively bounded, i.e.,
$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (26)
@@
Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent, by semantic measure satisfies additivity
$p_{i}(s,\theta)=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})-\mu_{C}(K_{0})-\sum_{j\neq i}\mu_{C}(K_{j})=\mu_{C}(K_{i}).$
Summing over all agents:
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$
Therefore, when $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we have $\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})$. By the independence assumption, $K_{i}\perp_{C}K_{j}\text{ for all }i\neq j$, and semantic information measure non-negativity, with $p_{i}=\mu_{C}(K_{i})\geq 0$,
$\mu_{C}(K_{i})=0\text{ for all }i=1,\ldots,n.$
Then, the nullset and positivity property of $\mu_{C}$ together with independence condition $K_{i}\perp_{C}K_{0}$ imply $K_{i}=\varnothing$ for all $i=1,\ldots,n$.
By the response generation principle:
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right)=\mathcal{E}_{C}(K_{0}\cup\varnothing)=\mathcal{E}_{C}(K_{0}).$
@@
$\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
@@
$K(g(s))\subseteq\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
@@
$\mu_{C}(K(g(s)))\leq\mu_{C}(B^{*}_{C}).$
@@
In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network. With CoT reasoning, multi-step explicit reasoning is iterative applications of reasoning rules within the computational budget $C$. In both cases, the emergence operator $\mathcal{E}_{C}$ bounds what can be derived. When information is conserved, even sophisticated CoT reasoning cannot create genuinely new information. It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$. This explains why CoT improves performance on complex reasoning tasks (by better utilizing $\mathcal{E}_{C}$) but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
Notice that in transformer architectures, information contribution $p_{i}$ can be estimated via attention head ablation:
$p_{i}\approx\tilde{J}(\text{model})-\tilde{J}(\text{model with head $i$ ablated}),$ (27)
where $\tilde{J}$ denotes the model training loss function.
@@ 5.5.3 Relevant Knowledge Revelation
Mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
@@ 5.5.4 Knowledge-Constrained Optimality
Mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\},$ (29)
i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only knowledge that is available and accessible through bounded reasoning.
@@ 6.1 Proof I: Inference with Independent Private Knowledge
Valuation of contributions depend on hallucination cost reduction:
$v_{j}(g(s),\theta_{j})=J(g((s_{-j},0_{j})),q)-J(g(s),q),$ (32)
and $J(r,q)$ strictly decreases when adding knowledge from $T(q)\setminus K(r)$.
Also, let us focus on non-trivial queries, integrating knowledge of at least two agents, so that:
$\mathcal{K}_{M}\cap T(q)\cap[K(\theta_{i})\triangle K(\theta_{j})]\neq\emptyset.$ (33)
and $(K(\theta_{i})\setminus K_{0})\cap(K(\theta_{j})\setminus K_{0})=\emptyset$ for $i\neq j$.
Theorem 6 (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries requiring independent agent-specific knowledge states $\theta$, no response mechanism $\mathcal{M}$ can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
@@
The proof goes by assuming the opposite. We assume all properties are satisfied, and show it leads to a contradiction.
By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, every truthful mechanism must use Groves transfers to measure individual information contributions:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (34)
where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy. Indeed, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
For relevant knowledge revelation, we need $u_{i}(s^{*}(\theta),\theta_{i})\geq 0$ for all $i$. The unique choice that ensures this while maintaining efficiency is Clarke pivotal rule:
$h_{i}(s_{-i})=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j}).$ (35)
That gives:
$p_{i}(s,\theta)$ $=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})=\Delta_{i}(s,\theta).$ (36)
Under knowledge-constrained optimality, the truthful strategy profile $s^{*}(\theta)$ minimizes hallucination cost by utilizing all relevant knowledge. Therefore, by strict monotonicity of $J(r,q)$ and disjoint extra knowledge, omitting any contributing agent $i$ strictly increases the hallucination cost:
$J(g((s_{-i}^{*},0_{i})),q)>J(g(s^{*}),q).$ (37)
Since agent’s valuation of contribution depends on $J(r,q)$, each agent $j\neq i$ faces higher loss when agent $i$ is removed, making their marginal contribution more valuable:
$v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})>v_{j}(g(s^{*}),\theta_{j}).$ (38)
Therefore, each Clarke pivot is strictly positive:
$\Delta_{i}(s^{*},\theta)=\sum_{j\neq i}[v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})-v_{j}(g(s^{*}),\theta_{j})]>0.$ (39)
Since each payment equals its Clarke pivot: $p_{i}(s^{*},\theta)=\Delta_{i}(s^{*},\theta)$, we have:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}(s^{*},\theta)>0.$ (40)
@@
That establishes a direct contradiction. Since Clarke pivotal payments are necessary for relevant knowledge revelation, and these necessarily sum to a positive value under knowledge-constrained optimality, the four properties cannot be simultaneously satisfied. ∎

### v5

@@ 4.1.2 The Uncertainty Resolution Paradox and its Resolution
Second, we prove that hallucinations are inevitable even when knowledge is subadditive (within a fixed context). In that setting, by Theorem 4, reasoning does not create new information. It is a transformation that reveals information already present in any knowledge representation. Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. However, we should take care of a paradox those claims apparently bring into existence.
@@
To resolve it, we need the emergence operator $\mathcal{E}_{C}(X|Q)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X|Q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time) in a given context. Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X|Q)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$. That is possible only with organized information providing knowledge for reasoning. These are precisely the statements of Theorem 1 and Theorem 4.
@@
$\theta(t+1)=\mathcal{E}_{C}(\theta(t)|Q)\quad\text{and}\quad y(t)=\mu_{C}(\theta(t)|Q)$ (14)
@@
$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)|Q).$ (15)
@@
$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$
$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$
$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$
We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)|Q)|Q)>\mu_{C}(\theta(0)|Q)$, i.e., practical information gain in the course of reasoning. In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$
@@ 4.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses by their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions). We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths. The following formal definition encapsulates those ideas.
@@
For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Indeed, that limit is given by the Goedel’s Incompleteness Theorems for the formalized mathematical theories. Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely. High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth. Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
@@
A query $q\in\mathcal{Q}$ is non-trivial with respect to computational budget $C$ if there exist knowledge configurations $\theta_{i}$ and $\theta_{j}$ for the context $Q=K(q)\cup T(q)$, such that
[…]
where $\triangle$ denotes symmetric difference.
A query is non-trivial when at least two agents disagree on some fact that is both encoded and available, and relevant to answering that query. For trivial queries impossibility results do not hold. That definition is important for the development of impossibility results. It measures the ability of a response generation system to store knowledge and expand knowledge that is necessary for the impossibility to emerge. We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework.
@@ 4.4 Hallucination Cost
If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$.
@@ 4.5.2 Semantic Information Conservation
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (22)
@@
Let $q\in\mathcal{Q}$ be a query with context $Q=K(q)\cup T(q)$ and let $B^{*}_{C}$ be the fixed point of reasoning for the baseline context-relevant knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$. Let us assume that the knowledge in the query-only response (without any inference-based contributions) meets the following condition:
@@
$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for all }i\neq j,$ (25)
$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{0}\cap\mathcal{R}_{C}(Q))\text{ for all }i=1,\cdots,n.$ (26)
@@
$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}|Q\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}|Q\bigr{)}.$ (27)
@@
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}|Q\right),$ (28)
@@
$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)|Q).$ (29)
If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$ and for all $f\in\mathcal{F}_{C}$ we have $f(A)\cap\mathcal{R}_{C}(Q)=\emptyset$ whenever $A\cap\mathcal{R}_{C}(Q)=\emptyset$, then the response is creatively bounded, i.e.,
$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (30)
@@
Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent within context $Q$, we have:
$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for }i\neq j.$
By the subadditivity axiom with equality for independent sets:
$p_{i}(s,\theta)=\mu_{C}(K_{0}|Q)+\sum_{j=1}^{n}\mu_{C}(K_{j}|Q)-\mu_{C}(K_{0}|Q)-\sum_{j\neq i}\mu_{C}(K_{j}|Q)=\mu_{C}(K_{i}|Q).$
Given $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we then have $\sum_{i=1}^{n}\mu_{C}(K_{i}|Q)=0$. By contextual non-negativity, $\mu_{C}(K_{i}|Q)\geq 0$, so we conclude that:
$\mu_{C}(K_{i}|Q)=0\text{ for all }i=1,\ldots,n.$
But by the contextual non-negativity axiom, $\mu_{C}(K_{i}|Q)=0$ if and only if $K_{i}\cap\mathcal{R}_{C}(Q)=\emptyset$. If follows that for every $f\in\mathcal{F}_{C}$ the irrelevance of partial knowledge is preserved under reasoning, i.e., $f(K_{i})\cap\mathcal{R}_{C}(Q)=\emptyset$. By the definition of the emergence operator:
$\mathcal{E}_{C}(K_{i}|Q)\cap\mathcal{R}_{C}(Q)=K_{i}\cap\mathcal{R}_{C}(Q)\cup\bigcup_{f\in\mathcal{F}_{C}}[f(K_{i})\cap\mathcal{R}_{C}(Q)]=\emptyset.$
By the chain-of-thought assumption and since emergence only produces relevant knowledge within $\mathcal{R}_{C}(Q)$, we get:
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\Big{|}Q\right)=\mathcal{E}_{C}(K_{0}|Q).$ (31)
@@
$\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
@@
$K(g(s))\subseteq\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
@@
$\mu_{C}(K(g(s))|Q)\leq\mu_{C}(B^{*}_{C}|Q).$
This completes the proof that semantic information conservation implies bounded creativity. ∎
@@
While computational independence (limiting synergetic creation of knowledge) is very restrictive, e.g., due to shared representations in neural networks, the theorem shows that with perfect separation of knowledge sources, semantic conservation prevents creative generation beyond baseline knowledge.
In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network. With CoT reasoning, multi-step explicit reasoning is iterative applications of reasoning rules within the computational budget $C$. In both cases, the emergence operator $\mathcal{E}_{C}$ bounds what can be derived. When information is conserved, even sophisticated CoT reasoning cannot create genuinely new information. It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$. This explains why CoT improves performance on complex reasoning tasks by better utilizing $\mathcal{E}_{C}$ but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
@@ 4.5.3 Relevant Knowledge Revelation
For query $q$ with context $Q=K(q)\cup T(q)$, mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$ with $K(\theta_{i})\cap\mathcal{R}_{C}(Q)\neq\emptyset$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
@@ 4.5.4 Knowledge-Constrained Optimality
For query $q$ with context $Q=K(q)\cup T(q)$, mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q)\cup K(q)|Q)\},$ (33)
i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only context-relevant knowledge accessible through bounded reasoning.
@@ 5.1 Proof I: Inference with Independent Private Knowledge
Valuation of contributions depend on hallucination cost reduction and convex $J(r,q)$ is strictly decreasing when adding relevant knowledge from $T(q)\cap\mathcal{R}_{C}(Q)\setminus K(r)$. Also, let us focus on non-trivial queries.
Theorem 6 (Impossibility Theorem for Independent Private Knowledge Settings).
For any query space $\mathcal{Q}$ containing non-trivial queries $q$ with context $Q=K(q)\cup T(q)$, where agents have independent knowledge states within $\mathcal{R}_{C}(Q)$, no response mechanism $\mathcal{M}$ can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
@@
Let us assume all properties are satisfied.
Consider $r=g(\theta)$ for any knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$. For a given context $q$ and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics. For the sake of clarity, consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big{(}K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big{)}$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
Also, for the fixed context let us encode the hallucination loss $J(r,q)$ as $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
$g(s)=g(s(\theta))\in\arg\max_{r\in\mathcal{R}}W(r,\theta)=\arg\max_{r\in\mathcal{R}}\sum_{i=1}^{n}v_{i}(r,\theta_{i})-v_{0}(r)\text{ for all }\theta.$ (36)
By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers to measure individual information contributions:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+v_{0}(g(s)),$ (37)
where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy. In particular, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule:
$h_{i}(s_{-i})=\max_{r\in\mathcal{R}}\Big{\{}\sum_{j\neq i}v_{j}(r,\theta_{j})-v_{0}(r)\Big{\}}.$ (38)
Notice that $g(\theta)$ maximizes $W(r,\theta)$, so for each $i$ we have:
$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big{(}g(\theta),\theta_{j}\big{)}-v_{0}\big{(}g(\theta)\big{)}=W(\theta)-v_{i}\big{(}g(\theta),\theta_{i}\big{)}.$ (39)
But that implies:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\Big{[}\sum_{j\neq i}v_{j}(g(s),\theta_{j})-v_{0}(g(s))\Big{]}\geq 0,$ (40)
and therefore:
$\sum_{i=1}^{n}p_{i}(s,\theta)\geq 0.$ (41)
By relevant knowledge revelation and non-triviality of $q$, there exist at least one pair of agents, say $i$ and $j$, for which
[…]
Then, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
$\sum_{i=1}^{n}p_{i}(s,\theta)>0.$ (42)
@@
That establishes a direct contradiction. The four properties cannot be simultaneously satisfied. ∎